A Level Mathematics 9709 Paper 62, May/June 2024

Question paper and mark scheme50 marks7 questions

Solve it on ExamPal

Practise

Untimed. Write on the paper and check as you go.

Start

Exam

Timed, with a clock you set, like sitting the real paper.

Start

Opens full screen.

Papers open in

Opening the question paper

Read as textThe questions and the mark scheme as plain text. Easier to read on a phone.

Question paper

9709/62 · May/June 2024

Download PDF
1

A random variable X has the distribution Po(145).

a

Use a suitable approximating distribution to calculate P(X⩽150).

[4]
b

Justify the use of your approximating distribution in this case.

[1]
2

Henri wants to choose a random sample from the 804 students at his college. He numbers the students from 1 to 804 and then uses random numbers generated by his calculator. The first 20 random digits produced by his calculator are as follows. 5 6 7 1 0 9 8 4 3 1 0 9 6 6 5 0 2 1 7 6 Henri’s first two student numbers are 567 and 109.

a

Use Henri’s digits to find the numbers of the next two students in the sample.

[2]
b

There were 30 students in Henri’s sample. He asked each of them how much time, X hours, they spent on social media each week, on average. He summarised the results as follows. n=30Σx=610Σx2=12405 Use this information to calculate an unbiased estimate of the mean of X and show that an unbiased estimate of the variance of X is less than 0.1 .

[3]
c

Henri’s friend claims that Henri has probably made a mistake in his calculation of Σx or Σx2. Use your answer to part (b) to comment on this claim.

[1]
3

A student wishes to estimate the proportion, p, of students at her college who have exactly one brother. She surveys a random sample of 50 students at her college and finds that 18 of them have exactly one brother. She calculates an approximate α% confidence interval for p and finds that the lower limit of the confidence interval is 0.244 correct to 3 significant figures. Find α correct to the nearest integer.

[4]
4

A random variable X has the distribution N(10,12). Two independent values of X, denoted by X1 and X2, are chosen at random.

a

Write down the value of P(X1>X2).

[1]
b

Find P(X1>2X2−3).

[5]
5

The number of goals scored by a sports team in the first half of any match has the distribution X∼Po(3.1). The number of goals scored by the same team in the second half of any match has the distribution Y∼Po(2.4). You may assume that the distributions of X and Y are independent.

a

Find P(X<4).

[2]
b

Find the probability that, in a randomly chosen match, the team scores at least 5 goals.

[3]
c

Given that the team scores a total of 5 goals in a randomly chosen match, find the probability that they score exactly 3 goals in the first half.

[4]
6

The masses of cereal boxes filled by a certain machine have mean 510 grams. An adjustment is made to the machine and an inspector wishes to test whether the mean mass of cereal boxes filled by the machine has decreased. After the adjustment is made, he chooses a random sample of 120 cereal boxes. The mean mass of these boxes is found to be 508 grams. Assume that the standard deviation of the masses is 10 grams.

a

Test at the 2.5% significance level whether the mean mass of cereal boxes filled by the machine has decreased.

[5]
b

Later the inspector carries out a similar test at the 2.5% significance level, using the same hypotheses and another 120 randomly chosen cereal boxes. Given that the mean mass is now actually 506 grams, find the probability of a Type II error.

[5]
7

The probability density function, f, of a random variable X is given by f(x)={k(1+cos⁡x)0⩽x⩽π,0otherwise, where k is a constant.

a

Show that k=1π.

[3]
b

Verify that the median of X lies between 0.83 and 0.84 .

[3]
c

Find the exact value of E(X).

[4]

Mark scheme

9709/62 · May/June 2024

Download PDF

1a · 4 marks

  • B1

    N(145, 145)

  • M1

    ±150.5−145145 [= ±0.457]

  • M1

    Φ(‘0.457’)

  • A1

    = 0.676 (3sf)

1b · 1 mark

  • B1

    145 > 15

2a · 2 marks

  • B2

    [567, 109], 665, 21

2b · 3 marks

  • B1

    Est(μ) =61030 or 613

  • M1

    Est(σ2) =3029(1240530−(61030)2) or 129(12405−610230)

  • A1

    = 0.0575 (3sf)

2c · 1 mark

  • B1

    Variance is [unrealistically] small so Henri has [probably] made a mistake/claim is [probably] correct

3 · 4 marks

  • M1

    1850−z×1850(1−1850)50=0.244

  • A1

    z=1.709 or 1.708

  • M1

    Φ−1(‘1.709’)=0.956 ; 1−2(1−‘0.956’) [= 0.912]

  • A1

    α=91

4a · 1 mark

  • B1

    0.5

4b · 5 marks

  • B1

    E(X1−2X2+3)=10−20+3 [= −7] or E(2X2−X1−3)=20−10−3 [= 7]

  • B1

    Var(X1−2X2+3)=12+22×12+0 [= 60]

  • M1

    0−(‘−7’)‘60’ [= 0.904]

  • M1

    1−Φ(‘0.904’)

  • A1

    = 0.183

5a · 2 marks

  • M1

    e−3.1(1+3.1+3.122!+3.133!) or e−3.1(1+3.1+4.805+4.965) or 0.0450+0.1397+0.2165+0.22368

  • A1

    = 0.625 (3sf)

5b · 3 marks

  • B1

    [λ]= 5.5

  • M1

    1−e−5.5(1+5.5+5.522!+5.533!+5.544!) or 1−e−5.5(1+5.5+15.125+27.7292+38.1276) or 1−e−5.5(0.004087+0.0224772+0.061812+0.113323+0.155819)

  • A1

    = 0.642 or 0.643 (3sf)

5c · 4 marks

  • M1

    [P(X=3)×P(Y=2)= ] = e−3.1×3.133!×e−2.4×2.422! or 0.223676×0.261267 [= 0.05844]

  • M1

    [P(total 5)= ] e−5.5×5.555! or 0.17140

  • M1

    P(P(exactly 3 in 1st half given total 5)=P(exactly 3 in 1st half and total 5)P(total 5)

  • A1

    [= ‘0.05844’‘0.17140’] = 0.341 (3sf)

6a · 5 marks

  • B1

    H0: Population mean mass = 510 g H1: Population mean mass < 510 g

  • M1

    ±508−51010÷120

  • A1

    = ±−2.191 or −2.190

  • M1

    −2.191<−1.96 or 2.191>1.96 Area comparison: 0.0143 or 0.0142 < 0.025

  • A1

    [Reject H0] There is sufficient evidence to suggest that the [mean] mass has decreased

6b · 5 marks

  • M1

    cv−51010÷120=−1.96

  • A1

    cv = 508.21

  • M1

    z=±508.21−50610÷120 [= 2.421]

  • M1

    P(X‾>508.21∣μ=506)=1−Φ(‘2.421’)

  • A1

    = 0.0077 to 0.0080 (2sf)

7a · 3 marks

  • M1

    k∫0π(1+cos⁡x)dx=1

  • A1

    k[x+sin⁡x]0π=1

  • A1

    [e.g. k(π+sin⁡π−(0+0))=1], kπ=1, k=1π

7b · 3 marks

  • M1

    1π[x+sin⁡x]00.83 or 1π(0.83+sin⁡0.83) 1π[x+sin⁡x]00.84 or 1π(0.84+sin⁡0.84)

  • A1

    = 0.499 (3 sf) = 0.504 (3 sf)

  • A1

    ‘0.499’ < 0.5 < ‘0.504’ hence 0.83 < median < 0.84 Equivalent to –0.000912 < 0 < 0.00441 hence 0.83 < median < 0.84

7c · 4 marks

  • M1

    1π∫0π(x+xcos⁡x)dx

  • 1 mark

    = 1π([x22]0π+[xsin⁡x]0π−∫0π(sin⁡xdx))

  • A1

    = 1π(x22+xsin⁡x+cos⁡x) or e.g. π2+1π(0−[−cos⁡x]0π) or e.g. π2+1π((−1−1))

  • A1

    = π2−2π

Questions

  1. 1
    Question 15 marks
    Mark scheme ›
  2. 2
    Question 26 marks
    Mark scheme ›
  3. 3
    Question 34 marks
    Mark scheme ›
  4. 4
    Question 46 marks
    Mark scheme ›
  5. 5
    Question 59 marks
    Mark scheme ›
  6. 6
    Question 610 marks
    Mark scheme ›
  7. 7
    Question 710 marks
    Mark scheme ›

More 9709 papers

Other papers from May/June 2024

Solve it on ExamPalStart